Q1-1An object has a net force LaTeX: \vec{F}=( F → = (14.9LaTeX: \hat{x}+x ^ +8.19LaTeX: \hat{y})Ny ^ ) N exerted on it. What is the magnitude of the acceleration (in unit of LaTeX: \frac{m}{s^2} m s 2) if the object has a mass of 5.46 kg?

Q2-An object has a net force LaTeX: \vec{F}=( F → = (24.2LaTeX: \hat{x}+x ^ +21.1LaTeX: \hat{y})Ny ^ ) N exerted on it. What is the magnitude of the force (in unit of newton)?
Answer #1

Q1)

magnitude of force, |F| = sqrt(14.9^2 + 8.19^2)

= 17 N

According to Newton's second law
magnitude of acceleration, a = |F|/m

= 17/5.46

= 3.11 m/s^2 <<<<<<----------Answer

Q2) magnitude of force, |F| = sqrt(24.2^2 + 21.1^2)

= 32.1 N <<<<<<----------Answer